Solve by completing the square.
x^2 + 14x + 48 = 0
My answer for this problem was -8, 6. Is this correct. Thanks so much for your time and help.

Answer by checkley75(3666)

X^2+14X+48=0
(X+8)(X+6)=0
X+8=0
X=-8 ANSWER.
X+6=0
X=-6 ANSWER.